a) Consider a silicon p-n junction at T = 300 K with doping concentrations of NA = 10^16 cm-3 . Given that the built-in voltage Vbi = 0.7 V, determine the doping concentration of the n-side of the junction


1
Expert's answer
2021-06-28T04:17:02-0400

Gives

T=300K

"N_A=10^{-16}cm^{-3}=10^{22}m^{-3}"

"N_n=?"

"V_{bi}=0.7V"

"N_n=N_Ae^\frac{V_{bi}e}{K_BT}"

Put value

"N_n=10^{22}e^\frac{0.7\times 1.6\times10^{-19}}{1.38\times10^{-23}\times300}"

"N_n=5.6\times10^{27}" Cm-3